OpenStudy (anonymous):

how to find the range of f(x)=2+sqrt(8x-2x^2)

OpenStudy (anonymous):

the square root of something is never negative, so the lower bound is \(y=2\)

OpenStudy (anonymous):

I want to see how to do it. My Teacher showed us different way of doing it

OpenStudy (anonymous):

to find the upper bound, find the maximum value of \(8x-x^2\)

OpenStudy (anonymous):

you know how to find that? it is the same as finding the second coordinate of the vertex

OpenStudy (anonymous):

can you show me?I want to make sure I am on the right track. please

OpenStudy (anonymous):

I got 2< or= f(x)<2+(sqrt.(8))

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{8}{2\times -2}=2\]

OpenStudy (anonymous):

second coordinate is what you get when you replace \(x\) by \(2\) namely 8 so your answer \[2\leq y\leq 2+\sqrt{8}\] is correct

OpenStudy (anonymous):

can you show me how will you get this one...because you may have an easier solution and easier to understand solution

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

your solution of how you will get this

OpenStudy (anonymous):

probably the same way you did it, since we got the same answer i wrote my method above find the second coordinate of the vertex of \(8x-2x^2\)

OpenStudy (anonymous):

this is how my professor did it: \[f(x)\ge2\] \[y=8x-2x ^{2}\] then he made a table of values... Is there an algebraic way of doing it?

OpenStudy (anonymous):

yes read the answer i wrote above

OpenStudy (anonymous):

can I see the solution please...I kinda get it

OpenStudy (anonymous):

i can write it again if you like, but i wrote it above the first coordinate of the vertex is \(8x-2x^2\) is \(-\frac{b}{2a}=-\frac{8}{2\times (-2)}=2\)

OpenStudy (anonymous):

the second coordinate of the vertex is \(8\times 2-2\times 2^2=8\)

OpenStudy (anonymous):

so the max of your function is \(2+\sqrt{8}\)

OpenStudy (anonymous):

Oh I get it...Thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

do I always have to use the equation under the root to find the vertex?

OpenStudy (anonymous):

depends on the function

OpenStudy (anonymous):

where did you get -b/2a?